1. WHICH ANGLE PRODUCES THE MAXIMUM HORIZONTAL DISTANCE?

2. HOW DOES CHANGING THE MASS AFFECT THE MOTION OF A PROJECTILE?

3. WHAT IS THE EFFECT OF AIR RESISTANCE?

4. HOW DOES CHANGING THE INITIAL SPEED AFFECT THE MOTION OF A PROJECTILE?
Answer #1

Answer:

(1) In projectile motion, horizontal distance R = v2Sin2\theta/g

R depends on the initial speed v and projectile angle \theta. Here, we will discuss only about angle dependence on the horizontal distance. If \theta = 0 then R = 0 and if \theta = 900 then again R = 0. So the best answer is if \theta = 450 then R will be maximum. Therefore, for maximum horizontal distance R we need to project the projectile with an angle \theta = 450.

(2) Here, we will know about the affect of mass on the projectile in its motion. In the projectile motion, mass does not affect the motion of a projectile. Mass does not affect either the velocity of the projectile or the horizontal distance. Mainly, motion of the projectile depends on the initial speed and angle and not on the mass m.

(3) In many projectile's motion problem we normally neglect the affect of air resistance on the projectile but in practical cases, air resistance can have a significant affect on a projectile's motion, especially if the motion lasts for some time. When the projectile is small and its speed of projection not too large then the air resistance can be described as a force opposing the motion and proportional to the velocity of the projectile. That means the air slows the projectile.

(4) Initial speed of a projectile is a very important parameter in the projectile's motion because of it can able to affect the maximum height of the projectile, maximum horizontal distance of the projectile and its time of flight.

Maximum height H = v2sin2\theta/2g

Horizontal distance R = v2sin2\theta/g

Time of flight t = 2vsin\theta/g

Here H, R and t all are the function of the initial speed and angle.